OpenStudy (anonymous):

find values of the trigonometric functions of the angle (in standard position) whose terminal side passes through the given points. give answer in exact form. (square root 3, 2)

OpenStudy (anonymous):

@Euler271 need help with this problem i did one similar did good on that but this one idk?

OpenStudy (anonymous):

for sin i get 1 and for cos i get 0 am i doing this right?

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ok i'm back

OpenStudy (wolf1728):

I drew a graphic. I'm guessing if the angle passes through the point (1.732, 2) then the angle would equal the arc tangent of 2/1.732 = 1.1547344111 = 49.107 degrees.

OpenStudy (wolf1728):

I changed the graphic to add the labels opp, adj and hyp.

OpenStudy (anonymous):

i've got that so far but not sure as to what to do. like round 1.73 to 4 for the radius and then do square root 2^2-2^2=0 ?

OpenStudy (wolf1728):

Well according to my graphic, it seems (to me) that you would have to solve for sine, cosine, tangent, cotangent, secant and cosecant of 49.107 degrees.

OpenStudy (anonymous):

i meant 4^2-2^2=12

OpenStudy (anonymous):

yes i would

OpenStudy (wolf1728):

Wouldn't the hypotenuse² equal 1.732² + 2²? And hypotenuse = sqrt (7) or about 2.646.

OpenStudy (anonymous):

the way you set it up looks right but the answer you have is not what i have but you could be right and i could be wrong

OpenStudy (wolf1728):

I'm guessing that you would have to find the value for all 6 trigonometric functions for 49.107 degrees ?

OpenStudy (wolf1728):

I'll be away from the keyboard for a while. (Taking my dog for a walk.)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I'll get back on here tomorrow and see what I come up with

OpenStudy (wolf1728):

Okay Ash90 - just send me an instant message and I'll come back too (unless of course you have figured it all out !)